OpenStudy (anonymous):

What is a possible first step to eliminate a variable in the following system? 2x – 3y = –1 3x – 4y = –3 Multiply the first equation by –3 and the second equation by 2. Multiply the first equation by 3 and the second equation by –2. Multiply the first equation by –4 and the second equation by 3. All of the above.

OpenStudy (anonymous):

@satellite73 could you help me on this?

OpenStudy (helder_edwin):

to eliminate a variable, it has to have the same coefficient but with different sign in each equation.

OpenStudy (anonymous):

alright whats next?

OpenStudy (helder_edwin):

u have ax+.. bx+.. how do u turn then numbers a and b into the same?

OpenStudy (anonymous):

i really dont know. im having trouble.

OpenStudy (helder_edwin):

*the numbers

OpenStudy (anonymous):

no i am having trouble :/

OpenStudy (helder_edwin):

u have ax+...=... bx+...=... if u multiply the first equation by b and the second by a. what do u get?

OpenStudy (anonymous):

i'm not sure what you are talking about

OpenStudy (anonymous):

i am terrible at math

OpenStudy (anonymous):

@satellite73 could you pleasehelp me

OpenStudy (helder_edwin):

say u have 3x-4y=2 2x-6y=3 and u multiply the first equation by 2 and the second by 3. so u get 6x-8y=4 6x-18y=9 what do u do next?

OpenStudy (helder_edwin):

well u change signs: say to the first equation, so u get -6x+8y=-4 6x-18y=9 now add both equation and get 0x-10y=5 or simply -10y=5 from which y=-1/2

OpenStudy (texaschic101):

I don't think she is online anymore

OpenStudy (helder_edwin):

now u know that y=-1/2 so replace this into either of your equations, say the second one: 2x-6y=3 becomes 2x-6(-1/2)=3 2x+3=3 2x=0 x=0 so the system is solved and x=0 and y=-1/2. try and do something similar with your exercise.